# Can Large Language Models Adequately Perform Symbolic Reasoning Over Time Series?

### v2

@@ Appendix B SymbolBench Dataset
    CDEs: We use an LLM to generate the skeleton of the continuous-time dynamical system, replacing each unknown coefficient with a placeholder "c". An optimizer then fits these coefficients to the training time series. To avoid the high cost of repeatedly solving and differentiating through a full ODE solver, we adopt the finite–difference approximation strategy from ODEFormer d’Ascoli et al. (2023). Once the functional form $\hat{f}(\mathbf{x};\boldsymbol{\phi})$ is obtained, where $\phi$ is the fitted coefficients, we generate the final numerical solution using SciPy’s scipy.integrate.solve_ivp, and compare the trajectory to the ground truth:
    The distribution of sample dims is shown in Figure 4(a).
@@
(b) Dimension distribution of Boolean Networks used in SymbolBench.
Appendix C LLMs for Benchmarking
We evaluate six representative LLMs, chosen for their diversity in size, training data, and specialization:
@@ Appendix C LLMs for Benchmarking
  4. (d)
  5. (e)
  6. (f)
Appendix D Baseline Implementation
  1. (a)
    CDEs: We follow ODEFormer’s baseline implementation and hyperparameter tuning d’Ascoli et al. (2023), using PySR with finite-difference approximations for skeleton search, and default greedy top-$k$ generation for ODEFormer.
  2. (b)
    BNs: We reimplement LogicGep’s Boolean-network inference using the same Geppy genetic-programming framework Gao et al. (2022), but omit the continuous-to-binary discretization and MLP-based constraint stages, since our training traces are already binary.
  3. (c)
    SCMs: All SCM baselines are based on Tigramite Runge et al. (2022), with the maximum time-lag set to 1 for fair comparison.
Appendix E Analysis on Table 2
@@
Appendix F Hybrid Method
We adopt two frameworks for hybrid approach by separately let make GP and LLMs play different roles. (a) As shown in Figure 5(a), in addition to the quantitative evaluation using MSE, an additional qualitative score produced by LLM is incorporated in the evolution loop; (b) As shown in Figure 5(b), in addition to the original closed loop (blue line), we provide an extended path (red line) that utilize GP to expand the history pool. From a different perspective, the initial population produced by LLMs also improves the generation for GP by providing context-enhanced initial populations.
@@ Appendix F Hybrid Method
Appendix G Convergence Rate
@@
Appendix H Example Outputs
H.1 Example predictions across three tasks
@@
H.2 Candidates of CoT reasoning output on coupled-differential equations
@@
H.3 RLM output for one sample with marked reasoning steps

### v3

@@ F.3 Error in Context Alignment
    x1 (v_Coup_fti): Transcription factor controlling cortical area identity
@@
    x3 (v_Fgf8): Growth factor signaling from anterior neural ridge
  4. (d)
    x4 (v_Pax6): Master regulator of cortical neurogenesis
  5. (e)
    x5 (v_Sp8): Transcription factor in frontal cortex specification
Analysis: The biological context constrains valid regulatory logic: Fgf8 and Sp8 typically act as upstream signaling molecules that inhibit posterior genes (Coup_fti, Emx2) while promoting anterior development. Pax6 functions as a competence factor that enables context-dependent responses to signaling gradients. This case demonstrates severe context misalignment with F1 = 0.087, showing fundamental violations of both computational and biological domain constraints:
@@
    Biological Hierarchy Inversion: The ground truth implements the established developmental cascade where anterior signals (x3: Fgf8, x5: Sp8) suppress posterior genes (x1: Coup_fti, x2: Emx2). However, the prediction creates feedback loops where x2 positively regulates x1, contradicting known developmental biology where these genes have antagonistic relationships.
@@
    Signaling Logic Violations: The ground truth correctly models mutual inhibition between anterior (x3, x5) and posterior (x1, x2) gene programs. The prediction fails to capture this essential developmental principle, instead creating cooperative interactions that would be biologically impossible during cortical regionalization.
The model’s symbolic search generated computationally invalid and biologically implausible structures, demonstrating how insufficient domain context can lead to fundamental misalignment with both mathematical and biological principles underlying Boolean network dynamics.
Appendix G Analysis on Table 2
@@
Appendix H Hybrid Method
We adopt two frameworks for hybrid approach by separately let make GP and LLMs play different roles. (a) As shown in Figure 8(a), in addition to the quantitative evaluation using MSE, an additional qualitative score produced by LLM is incorporated in the evolution loop; (b) As shown in Figure 8(b), in addition to the original closed loop (blue line), we provide an extended path (brown line) that utilize GP to expand the history pool. From a different perspective, the initial population produced by LLMs also improves the generation for GP by providing context-enhanced initial populations.
@@ Appendix H Hybrid Method
Appendix I Convergence Rate
@@
Appendix J Example Outputs
J.1 Example predictions across three tasks
@@
J.2 Candidates of CoT reasoning output on coupled-differential equations
@@
J.3 RLM output for one sample with marked reasoning steps